# Learning Force Distribution Estimation for the GelSight Mini Optical Tactile Sensor Based on Finite Element Analysis

### v1

@@ Abstract
Contact-rich manipulation remains a major challenge in robotics. Optical tactile sensors like GelSight Mini offer a low-cost solution for contact sensing by capturing soft-body deformations of the silicone gel. However, accurately inferring shear and normal force distributions from these gel deformations has yet to be fully addressed. In this work, we propose a machine learning approach using a U-net architecture to predict force distributions directly from the sensor’s raw images. Our model, trained on force distributions inferred from Finite Element Analysis (FEA), demonstrates promising accuracy in predicting normal and shear force distributions. It also shows potential for generalization across sensors of the same type and for enabling real-time application. The codebase, dataset and models are open-sourced and available at https://feats-ai.github.io.
@@ I Introduction
Figure 1: Complete Method Overview: from data collection to force distribution prediction. After data collection in a precisely calibrated setup with a CNC milling machine, Finite Element Analysis is employed to generate labels (“ground truth” force distributions). Using the labels and raw images captured by the GelSight Mini tactile sensor, we train a U-net for efficiently mapping raw tactile images to the corresponding force distributions.
Tactile sensing plays an important role in advancing the state-of-the-art in robotic manipulation [1, 2, 3, 4, 5, 6, 7, 8]. Successful applications include grip adaptation through slip detection [9, 10, 11], medical procedures [12, 13] and tele-operation [14].
In particular, optical tactile sensors have emerged as a promising technology for capturing contact information due to their high spatial resolution, multimodal sensing capabilities—including shape [15], hardness [16], texture [17], and temperature [18]—and cost-effectiveness [19, 13]. However, many prior works have focused on extracting only low-dimensional tactile information, such as total force [20, 21, 22], limiting operational flexibility. Access to contact force distributions, on the other hand, would enable better handling of multiple contacts and diverse manipulation scenarios [23].
@@
Recent works leverage Deep Learning to address the challenge of real-time force estimation. In [28], Convolutional Neural Networks (CNNs) were used to predict contact forces from sensor images, while [29] introduced CANFnet for estimating normal force distributions at the pixel level. In [23, 30, 31], FEA-derived data was used to train a model for predicting force distributions, demonstrating the effectiveness of combining simulations with data-driven methods.
In this paper we introduce FEATS (see Fig. 1)—a machine learning approach that directly maps raw tactile images to force distributions, building upon the method by Sferrazza et al. [31]. We utilize FEA to generate labeled data for training, ensuring accurate ground truth across various indenters and force levels. A U-net neural network architecture [32] is employed to estimate force distributions from images captured by the GelSight Mini optical sensor [28, 33]. In contrast to [31], our method is tailored to a widely available commercial sensor GelSight Mini, dropping the requirement of a custom-made gel with immersed particles, thereby drastically extending the applicability of the approach. Furthermore, this sensor allows for a significantly expanded range of measurable forces $0-40\,\mathrm{N}$, an $8$-fold increase in the maximum measurable force compared to [31]. Finally, we open source our code, dataset and model, aiding reuse and reproducibility.
Experimental results demonstrate that the proposed method accurately predicts high-dimensional contact force distributions from raw tactile images. This capability advances robotic manipulation by accommodating a wider range of contact scenarios and offers a versatile representation applicable to downstream tasks.
@@ II-C Force Distribution Prediction Through Elasticity Theory
Figure 2: Models of 3D-printed indenters used for data collection. Different colors represent groups of indenters with similar shapes.
@@
However, direct prediction of force distributions from displacements, usually through a linear stiffness matrix, does not account for the nonlinearities of soft elastomers. Sun et al. [40] addressed this limitation by employing ResNet [41], which was trained on sensor images with approximated force distributions. Similarly, Sferrazza et al. [31] utilized a Deep Neural Network (DNN) trained on image features with force distributions obtained from FEA, and later improved this method by incorporating simulated training data [23, 30].
@@ III Method
Figure 3: Label creation process when a spherical indenter presses into the sensor’s soft silicone gel. The visualized contact force distributions correspond to the shear force’s $x$ component. Left: Simulation of the contact configuration and the raw output from running the 3D FEA. Middle: Projecting the result from the 3D FEA into the coordinate system of Gelsight Mini, i.e., into an image plane. Right: Force labels after changing the resolution to $24\times 32$.
Our proposed method—Finite Element Analysis for Tactile Sensing (FEATS)—estimates the force distribution acting upon the gel of the GelSight Mini sensor by approximating the output of a Finite Element Analysis (FEA) computation with a neural network that takes the raw RGB image as input. Querying a neural network is much faster than running FEA, and importantly, for FEA computation, one needs a precise geometrical description of the contact, whereas FEATS only needs a raw input image and no further geometric information. Thus, FEATS enjoys fast inference time and does not require any additional object tracking equipment at run time. Only at training time FEATS requires a dataset collected in a controlled environment on a set of calibration objects.
@@ III-A Data Collection
The data collection process involves a series of precise indentation experiments. We attach a GelSight Mini sensor to the fixed spindle of a Computer Numerical Control (CNC) milling machine, with a positional tolerance of ±$\pm 0.25\,\mathrm{\mu m}$ to automatically create different contact configurations between the sensor and selected indenters (cf. Fig. 1). Following [30], we use $12$ indenters of varying shapes and sizes (see Fig. 2). The Gelsight Mini sensor is equipped with the dotted gel to better track the indentation motion. We additionally place a six-axis RESENSE-HEX-21 Force/Torque (F/T) sensor above the Gelsight Mini to have a complementary external force measurement, which is later used to validate the material model. With this setup, we collected a total of $5173$ samples. Each sample contains the respective GelSight Mini RGB image, the CNC motion data, i.e., the current indentation, and the F/T sensor’s readings. The forces in the $z$ direction reach up to $40\,\mathrm{N}$, and in the $x$ and $y$ directions up to ±$\pm 5\,\mathrm{N}$.
@@ III-B Finite Element Analysis
For simulating the indentation experiments, we employ FEA using the open-source solver CalculiX [42]. The solver’s capability for nonlinear computations makes it an appropriate tool for the stress analysis of soft elastomers. Consequently, it can calculate the resulting force distributions corresponding to the previously described real-world indentation experiments – an essential component for label generation as described in the next Section III-C. We generate tetrahedral volume meshes for both the gel and indenters, with denser meshing at the contact surface to ensure higher accuracy of the simulated contact forces (see Fig. 3). The gel mesh comprises $2504$ contact surface elements out of a total of $7591$ elements. The elements are implemented as ten-node tetrahedral (C3D10 [42]) elements. The FEA is conducted as a static analysis, assuming hard contact and applying tied contact constraints. This approach permits the execution of simulations without requiring a friction coefficient, and is substantiated by the fact that the experiments are conducted in a manner that precluded slippage. Assuming no deformation of the indenters, they are characterized as a hard material with a Young’s modulus of 210 GPa and a Poisson ratio of 0.3. In [28], the gel elastomer of the GelSight sensor is characterized as a material similar to a Neo-Hookean solid with a shear modulus, $\mu$, of $0.145$. When employing a hyperelastic Neo-Hookean model in the CalculiX solver [42], the strain energy potential is expressed as $U=C_{10}(\bar{I}_{1}-3)+1/D_{1}(J-1)^{2}$ where $\bar{I}_{1}$ denotes the first invariant of the right Cauchy-Green deformation tensor, while $J$ represents the determinant of the deformation gradient tensor. $C_{10}$ and $D_{1}$ are the material constants to be set. In agreement with [28], $C_{10}$ is chosen to be $0.0725$. Due to CalculiX’s lack of support for perfectly incompressible materials, the solver assigns a default value to $D_{1}$ prior to the simulation. This model is later validated in the results section (cf. Sec. IV-A) by optimizing the material parameters through load-depth indentation data [43], comparing the normal forces from the FEA with F/T sensor measurements.
@@ III-C Creating Labels
Once all nodes have been projected onto the image plane, the force distributions for shear and normal forces are binned within the image boundaries. The number of bins decides the resolution of the force distribution. In most experiments, it is $24\times 32$, however, the resolution can be seamlessly adapted. Each bin’s force value is calculated by summing the contact force contributions of all elements intersecting that bin (see Fig. 3), where the element position refers to its state before deformation. The contribution of each element to a bin is proportional to the fraction of the element that falls within the bin.
@@ III-D U-net for Learning Force Distribution Estimation
The dataset is split into three parts: $85\%$ are used for training, $5\%$ for validation, and $10\%$ for testing. Prior to training, both input images and the ground truth force distribution labels are normalized with respect to the training dataset to ensure a consistent baseline across the data. Shear forces are normalized to a range of $-1$ to $+1$, while normal forces range from 00 to $+1$. Data augmentation techniques, such as adding Gaussian noise and adjusting image brightness, contrast, saturation, and hue, are used to enhance the model’s generalization capabilities. Training is carried out using the Adam optimizer [45], with an initial learning rate of $0.001$ and a batch size of $8$. The learning rate is adjusted adaptively based on validation performance. As validation loss, the Mean Absolute Error (MAE) on the sum of total forces in the unnormalized space is used. We train for $600$ epochs and select the model with the lowest validation loss.
@@ IV-A Material Characterization
To validate the Neo-Hookean material model parameter $C_{10}$, load-depth indentation data can be utilized (cf. Sec. III-B). The load-depth measurements are obtained, using a sphere indenter with a $15\,\mathrm{mm}$ diameter, by sampling data points at different indentation depths ranging from 0.5 mm to 2.0 mm, increasing in 0.5 mm intervals. The measured normal forces from the F/T sensor $\tilde{f}^{(i)}_{z}$ are compared with the forces $f^{(i)}_{z}(C_{10})$ calculated by the FEA via the MSE loss
@@
Figure 5: Ground truth labels (left column) and predictions of our FEATS model (right column) of the force distributions, in Newtons. The sphere indenter (cf. Fig. 3) penetrates the gel for $1.2\,\mathrm{mm}$, exerting a significant normal force. The resulting gel deformation causes shear forces, in this case roughly cancelling each other due to the absence of horizontal movement of the indenter.
@@ IV-B Evaluation of U-net Force Distribution Estimation
Table I: U-net Mean Absolute Error (MAE) on the Test Set
@@
$f_{x}$ $0.0006\pm 0.0006$ $0.2242\pm 0.4007$
$f_{y}$ $0.0005\pm 0.0003$ $0.0934\pm 0.1356$
$f_{z}$ $0.0015\pm 0.0010$ $0.3720\pm 0.4727$
To assess the U-net’s capacity to accurately estimate shear and normal force distributions from the raw RGB images across a wide range of forces, after training, we use a test dataset consisting of $492$ samples. The test data is recorded using the same sensor and approach as for the training data, described in Sec. III-A. The value ranges of the ground truth force distribution in the test dataset for the Total Force (TF) estimation task are shown in Fig. 6. Given the nature of the indentation experiments, the GelSight Mini sensor primarily encounters normal forces, while the range of the encountered shear forces is comparatively smaller.
Evaluation of U-net Predictions on the Test Dataset
@@
Our main U-net model outputs the force distribution with dimensions $24\times 32\times 3$, striking a good balance between maintaining a sufficiently high spatial resolution and at the same time low error on the test data (cf. Tables I and II). An example of the model prediction is shown in Fig. 5, resulting from the contact with a spherical object as in Fig. 3. The U-net demonstrates a noteworthy capability to predict shear and normal force distributions which closely align with the ground truth data (cf. left vs. right column in Fig. 5). This visual comparison highlights the U-net’s effectiveness in capturing the fundamental structures of the underlying force distribution patterns. Additional examples that underscore this finding are provided on the accompanying website.
@@
In order to quantitatively evaluate the performance of predicting force distributions, the MAE metric was used. With MAE, the deviations between the predicted and ground truth force values can be measured, both for per-Grid Unit Forces (GUF) and for the Total Force (TF) values. We are showing the GUF errors since they directly reflect the error in the model outputs and in addition capture the contact area coverage, whereas TF is obtained as the sum of the GUFs and therefore is missing the local force information. The results of the main U-net evaluation on the test dataset are shown in Table I. The U-net is capable of accurately predicting the total shear and normal forces, with MAE values bellow $1\,\mathrm{N}$. It has been shown that the U-net performs better when dealing with shear forces in the $x$- and $y$-directions than with normal forces in the $z$-direction. This difference in performance may be attributed to two factors. First, the range of force values encountered by the U-net for shear forces are generally smaller than for normal forces. Second, the markers in the GelSight mini sensor’s gel inherently encompass a richer set of features for extracting and representing shear behavior. Shear forces cause a greater displacement of the markers compared to normal forces, and might therefore trigger a clearer signal.
Impact of Label Resolution on Prediction Accuracy
We compare our main U-net model with variants of the U-net architecture and against a ResNet to investigate whether changing the output resolution and network structure affects the prediction quality. The results on the Total Force (TF) estimation task in Table II show that the ResNet, which solely regresses to the total force, outperforms all considered U-net architectures in predicting the shear forces but is significantly less accurate in predicting the normal forces. Importantly, the ResNet only outputs a single 3D vector of force for the whole sensor, and therefore significantly lacks in resolution.
The top performing model according to the MAE on the total force is the U-net with output dimension of $12\times 16\times 3$. Within the variance of the results, it is however very close to our main architecture with output $24\times 32\times 3$, outperforming it on $x$-shear force but being worse on the $z$-normal force. That is why we recommend the network with a higher-resolution output, as it strikes the best balance between resolution and achieving small error on the test dataset.
We also compare against training three separate U-net models, where each model is designed to solely predict a single force component, i.e., each individual model outputs a force distribution of shape $24\times 32\times 1$. Although these models collectively possess three times the number of parameters, their predictive accuracy is generally inferior to that of the main U-net. This suggests that predicting all three force directions simultaneously can capture correlations between normal and shear forces, yielding more accurate predictions.
Lastly, when training a U-net model with shape $48\times 64\times 3$, its predictive accuracy remains relatively strong but exhibits a slight decline compared to our main U-net.
Generalization to Different Sensors
Table II: Model Ablation on the Total Force Estimation Task
@@
$\text{ResNet}^{\text{\tiny$1{\times}3$}}$ $\mathbf{0.085\pm 0.115}$ $\mathbf{0.069\pm 0.085}$ $1.593\pm 1.131$
@@
To explore the generalization capability of our FEATS method, it is tested on images captured with a different GelSight Mini sensor of the same type (cf. Fig. 7). The predictions remain largely accurate, although there is a notable downward and slight rightward shift in the localization of the contact area. This shift is likely due to differences in the alignment of the camera relative to the gel, thus causing the markers in the image of the test sensor to appear lower than in the image of the GelSight Mini sensor used during training. Furthermore, sensor-related factors such as image brightness, contrast, saturation and hue may also contribute to the shift in the localization of the contact area.
@@
Figure 7: Generalization to a different sensor. We evaluate our FEATS U-net on indenters from the test set using a different GelSight Mini sensor to validate the model generalization capability. The first row displays the raw images captured by the sensor, and the subsequent rows show the components of the force distribution, in Newtons. While the general outline of the force distribution is captured correctly, there is a noticeable downward and slight rightward shift in the localization of the predicted forces, likely due to different alignments of the gels in different sensors.

### v2

@@ Abstract
Contact-rich manipulation remains a major challenge in robotics. Optical tactile sensors like GelSight Mini offer a low-cost solution for contact sensing by capturing soft-body deformations of the silicone gel. However, accurately inferring shear and normal force distributions from these gel deformations has yet to be fully addressed. In this work, we propose a machine learning approach using a U-net architecture to predict force distributions directly from the sensor’s raw images. Our model, trained on force distributions inferred from Finite Element Analysis (FEA), demonstrates promising accuracy in predicting normal and shear force distributions for the commercially available GelSight Mini sensor. It also shows potential for generalization across indenters, sensors of the same type, and for enabling real-time application. The codebase, dataset and models are open-sourced and available at https://feats-ai.github.io.
@@ I Introduction
Figure 1: Method Overview from data collection to force distribution prediction. Starting from collecting versatile indentation data in a precisely calibrated setup with a CNC milling machine, we employ Finite Element Analysis for label generation, i.e., obtaining the corresponding “ground truth” force distributions. Through the combination of labels and the raw tactile images, we train a U-net capable of efficiently mapping raw tactile images to the corresponding force distributions during inference.
Tactile sensing plays an important role in advancing the state-of-the-art in robotic manipulation [1, 2, 3, 4, 5, 6, 7, 8]. Successful applications include grip adaptation through slip detection [9, 10, 11, 12], medical procedures [13, 14] and tele-operation [15].
In particular, optical tactile sensors have emerged as a promising technology for capturing contact information due to their high spatial resolution, multimodal sensing capabilities [16]—including shape [17], hardness [18], texture [19], and temperature [20]—and cost-effectiveness [21, 14]. However, many prior works have focused on extracting only low-dimensional tactile information, such as total force [22, 23, 24], limiting operational flexibility. Access to contact force distributions, on the other hand, would enable better handling of multiple contacts and diverse manipulation scenarios [25].
@@
Recent works leverage Deep Learning to address the challenge of real-time force estimation. In [30], Convolutional Neural Networks (CNNs) were used to predict contact forces from sensor images, while [31] introduced CANFnet for estimating normal force distributions at the pixel level. In [25, 32, 33, 34], FEA-derived data was used to train a model for predicting force distributions, demonstrating the effectiveness of combining simulations with data-driven methods. Yet, [33, 25, 32, 34] rely on custom-made optical tactile sensors, thereby imposing a significantly increased entry barrier into the field as the models are not compatible with off-the-shelf, commercially available sensors.
Inspired by these works, this paper introduces FEATS (see Fig. 1)—a machine learning approach directly mapping raw tactile images to force distributions. FEATS is tailored to the commercially available GelSight Mini optical tactile sensor [30, 35], thereby dropping the need for sensor manufacturing and drastically extending the applicability of the approach. Furthermore, this sensor allows for a significantly expanded range of measurable forces $0-40\,\mathrm{N}$, an $8$-fold increase in the maximum measurable force compared to [33], and an $10$-fold increase compared to [34]. FEATS is trained through an extensive dataset of real-world indentation data and the corresponding force labels obtained from the open-source FEA solver CalculiX [36]. To ensure high data quality, we collect the data using a CNC milling machine. On the representation level, FEATS leverages an efficient U-net neural network architecture [37], which is trained to map from a single GelSight image to the corresponding force distribution during inference. We optimize the network architecture for providing accurate force predictions while maintaining low inference times suitable for real-time control. Our thorough experimental results demonstrate that FEATS accurately predicts high-dimensional contact force distributions from raw tactile images and generalizes to novel, previously unseen indenters. Moreover, we showcase that the underlying U-Net architecture outperforms several baselines varying w.r.t. network size and output resolution.
To summarize, our key contributions are i) a method for creating force labels from FEA outputs tailored to the commercially available GelSight Mini, ii) the collection of a large dataset of versatile indentation data, iii) a model architecture capable of generalizing across objects and allowing for real-time inference, and iv) open-sourcing our dataset, the code for data collection, label generation, and model training, as well as the trained models for aiding reuse and reproducibility.
@@ II-C Force Distribution Prediction Through Elasticity Theory
Figure 2: Experimental setup for data collection. We use a CNC milling machine to create different calibrated contact configurations between the GelSight Mini tactile sensor and 3D printed indenters. We also place a six-axis F/T sensor above the tactile sensor to obtain complementary external force measurements.
@@
However, direct prediction of force distributions from displacements, usually through a linear stiffness matrix, does not account for the nonlinearities of soft elastomers. Sun et al. [44] addressed this limitation by employing ResNet [45], which was trained on sensor images with approximated force distributions. Similarly, Sferrazza et al. [33] utilized a Deep Neural Network (DNN) trained on image features with force distributions obtained from FEA, and later improved this method by incorporating simulated training data [25, 32]. A similar approach was also adopted in [34], where the authors additionally investigated soft objects interacting with the tactile sensor.
@@ III Method
Our proposed method—Finite Element Analysis for Tactile Sensing (FEATS)—estimates the force distribution acting upon the gel of the GelSight Mini sensor by approximating the output of a Finite Element Analysis (FEA) computation with a neural network that takes the raw RGB image as input. Querying a neural network is much faster than running FEA, and importantly, for FEA computation, one needs a precise geometrical description of the contact, whereas FEATS only needs a raw input image and no further geometric information. Thus, FEATS exhibits low inference times and does not require any additional object tracking equipment at run time. Only at training time, FEATS requires a labeled dataset collected in a controlled environment on a set of calibration objects, so-called indenters (cf. Fig. 2).
@@ III-A Data Collection
The data collection process involves a series of precise indentation experiments. We attach a GelSight Mini sensor to the fixed spindle of a Computer Numerical Control (CNC) milling machine, with a positional tolerance of ±$\pm 0.25\,\mathrm{\mu m}$ to automatically create random contact configurations between the sensor and selected indenters (cf. Fig. 2). This work leverages $24$ 3-d printed indenters ($19$ for training and $5$ for testing) varying in shape and size as shown at the bottom of Fig. 2. The optical tactile sensor, i.e., Gelsight Mini, is equipped with the dotted gel to better track the indentation motion. We additionally place a six-axis RESENSE-HEX-21 Force/Torque (F/T) sensor above the Gelsight Mini to have a complementary external force measurement, which is later used to validate the material model. With this setup, we collected a total of $14279$ samples. Each sample contains the respective GelSight Mini RGB image, the CNC motion data, i.e., the current indentation, and the F/T sensor’s readings. The forces in the $z$ direction reach up to $40\,\mathrm{N}$, and in the $x$ and $y$ directions up to ±$\pm 5\,\mathrm{N}$.
Figure 3: Label creation process when a spherical indenter presses into the sensor’s soft silicone gel. From left to right: 1) GelSight Mini image, 2) Simulation of the contact configuration and the raw output from running the 3D FEA, 3) Projecting the result from the 3D FEA into the coordinate system of Gelsight Mini, i.e., into an image plane, 4) Force labels after changing the resolution to $24\times 32$.
@@ III-B Finite Element Analysis
For simulating the indentation experiments, we employ Finite Element Analysis (FEA) using the open-source solver CalculiX [36]. The solver’s capability for nonlinear computations makes it an appropriate tool for the stress analysis of soft elastomers. Consequently, it can calculate the resulting force distributions corresponding to the previously described real-world indentation experiments – an essential component for label generation as described in the next Section III-C. We generate tetrahedral volume meshes for both the gel and indenters, with denser meshing at the contact surface to ensure higher accuracy of the simulated contact forces (see Fig. 3). The gel mesh comprises $2504$ contact surface elements out of a total of $7591$ elements. The elements are implemented as ten-node tetrahedral (C3D10 [36]) elements. The FEA is conducted as a static analysis, assuming hard contact and applying tied contact constraints. This approach permits the execution of simulations without requiring a friction coefficient, and is substantiated by the fact that the experiments are conducted in a manner that precluded slippage. Assuming no deformation of the indenters, they are characterized as a hard material with a Young’s modulus of 210 GPa and a Poisson ratio of 0.3. In [30], the gel elastomer of the GelSight sensor is characterized as a material similar to a Neo-Hookean solid with a shear modulus, $\mu$, of $0.145$. When employing a hyperelastic Neo-Hookean model in the CalculiX solver [36], the strain energy potential is expressed as $U=C_{10}(\bar{I}_{1}-3)+1/D_{1}(J-1)^{2}$ where $\bar{I}_{1}$ denotes the first invariant of the right Cauchy-Green deformation tensor, while $J$ represents the determinant of the deformation gradient tensor. $C_{10}$ and $D_{1}$ are the material constants to be set. In agreement with [30], $C_{10}$ is chosen to be $0.0725$. Due to CalculiX’s lack of support for perfectly incompressible materials, the solver assigns a default value to $D_{1}$ prior to the simulation. This model is later validated in the results section (cf. Sec. IV-A) by optimizing the material parameters through load-depth indentation data [46], comparing the normal forces from the FEA with F/T sensor measurements.
@@ III-C Creating Labels
Once all nodes have been projected onto the image plane, the force distributions for shear and normal forces are binned within the image boundaries. The number of bins decides the resolution of the force distribution. In most experiments, it is $24\times 32$, however, the resolution can be seamlessly adapted. Each bin’s force value is calculated by summing the contact force contributions of all elements intersecting that bin (see Fig. 3), where the element position refers to its state before deformation. The contribution of each element to a bin is proportional to the fraction of the element that falls within the bin. This process is repeated for each force component. Note that all the code for label generation is open-sourced on our website.
@@ III-D U-Net for Learning Force Distribution Estimation
We train the model using $9084$ samples and use a validation dataset containing $538$ samples. We later evaluate the model on three different test datasets for which we provide the details in the respective experimental sections. Prior to training, both input images and the ground truth force distribution labels are normalized with respect to the training dataset to ensure a consistent baseline across the data. Shear forces are normalized to a range of $-1$ to $+1$, while normal forces range from 00 to $+1$. Data augmentation techniques, such as adding Gaussian noise and adjusting image brightness, contrast, saturation, and hue, are used to enhance the model’s generalization capabilities. Training is carried out using the Adam optimizer [47], with an initial learning rate of $0.001$ and a batch size of $8$. The learning rate is adjusted adaptively based on validation performance. As validation loss, the Mean Absolute Error (MAE) on the sum of total forces in the unnormalized space is used. We train for $400$ epochs and select the model with the lowest validation loss.
@@ IV-A Material Characterization
To validate the Neo-Hookean material model parameter $C_{10}$, load-depth indentation data can be utilized (cf. Sec. III-B). The load-depth measurements are obtained, using a sphere indenter with a $15\,\mathrm{mm}$ diameter, by sampling data points at different indentation depths ranging from 0.5 mm to 2.0 mm, increasing in 0.5 mm intervals. The measured normal forces from the F/T sensor in our experimental setup $\tilde{f}^{(i)}_{z}$ are compared with the forces $f^{(i)}_{z}(C_{10})$ calculated by the FEA via the MSE loss
@@
Figure 5: Ground truth labels (left column) and predictions of our FEATS model (right column) of the force distributions, in Newtons. The sloping cuboid indenter (cf. Fig. 2) penetrates the gel for $1.23\,\mathrm{mm}$, exerting a significant normal force. The resulting gel deformation causes shear forces, in this case roughly cancelling each other due to the absence of horizontal movement of the indenter.
@@ IV-B Evaluation of U-net Force Distribution Estimation
Next, we assess the U-nets’ capacity to accurately estimate shear and normal force distributions from the raw RGB images of a GelSight Mini sensor. We evaluate the trained networks on multiple test datasets that cover a wide range of forces and contact configurations. The test dataset collection follows the same procedure as described in Secs. III-A & III-C).
The following experimental evaluation is split into four parts and aims to answer the following questions: (a) How well does the method perform on a test dataset containing different indentations with the same indenters as seen during training?, (b) How does the network architecture and output resolution affect the force prediction accuracy?, (c) Can the network generalize to novel, previously unseen testing indenters?, and (d) Can the network generalize to a different GelSight Mini sensor?
Throughout all the experiments, to quantitatively evaluate the performance of predicting force distributions, we employ the MAE metric. With MAE, the deviations between the predicted and ground truth force values can be measured, both for per-Grid Unit Forces (GUF) and for the Total Force (TF) values. We are reporting the GUF errors in addition to the TF errors since they directly reflect the error in the models’ individual output values. They, therefore, also capture spatial information regarding the contact area, whereas TF is obtained as the sum of the GUFs and is therefore missing out on the local information.
Evaluation of U-Net Predictions on the Test Dataset
Table I: U-net Mean Absolute Error (MAE) on the Test Set containing known Indenters.
@@
$f_{x}$ $0.0010\pm 0.0013$ $0.2932\pm 0.4173$
$f_{y}$ $0.0010\pm 0.0012$ $0.2302\pm 0.3405$
$f_{z}$ $0.0030\pm 0.0029$ $0.7994\pm 1.0144$
@@
This first evaluation considers our main U-net model that is trained to output the force distribution with dimensions $24\times 32\times 3$, therefore striking a good balance between maintaining a sufficiently high spatial resolution and at the same time low error on the test data. The test dataset contains $1068$ samples. This dataset is recorded using the same sensor and the same indenters as used during training data collection, i.e., the train indenters (cf. Fig. 2). The value ranges of the ground truth force distribution in the test dataset for the Total Force (TF) estimation task are shown in Fig. 6. Given the nature of the indentation experiments, the GelSight Mini sensor therefore primarily encounters normal forces, while the range of the encountered shear forces is smaller.
@@
An example of the model prediction is shown in Fig. 5, resulting from the contact with a sloping cuboid indenter. The U-net demonstrates a noteworthy capability to predict shear and normal force distributions which closely align with the ground truth data (cf. left vs. right column in Fig. 5). This visual comparison highlights the U-net’s effectiveness in capturing the fundamental structures of the underlying force distribution patterns. Additional examples that underscore this finding are provided on the accompanying website.
@@
Quantitative results are also provided in Table I. The U-net is capable of accurately predicting the total shear and normal forces, with mean MAE values below $1\,\mathrm{N}$. It has been shown that the U-net performs better when dealing with shear forces in the $x$- and $y$-directions than with normal forces in the $z$-direction. This difference in performance may be attributed to two factors. First, the range of force values encountered by the U-net for shear forces are generally smaller than for normal forces. Second, the markers in the GelSight mini sensor’s gel inherently encompass a richer set of features for extracting and representing shear behavior. Shear forces cause a greater displacement of the markers compared to normal forces, and might therefore trigger a clearer signal.
Table II: Model Ablation on the Total Force Estimation Task. Colors indicate the three best ranked methods - blue indicates the best, orange the second best and violet the third best.
@@
[…]
Impact of Label Resolution on Prediction Accuracy
This section compares our main U-net model with variants of the U-net architecture and against a ResNet to investigate whether changing the output resolution and network structure affects the prediction quality. We consider the same testing dataset as in the previous experiment.
The results on the Total Force (TF) estimation task in Table II show that the ResNet, which solely regresses to the total force, outperforms all considered U-net architectures in predicting the shear forces but is significantly less accurate in predicting the normal forces. Importantly, the ResNet only outputs a single 3D vector of force for the whole sensor, and therefore significantly lacks spatial resolution. The top performing model according to the MAE on the total force is the U-net with output dimension of $12\times 16\times 3$. Within the variance of the results, it is however very close to our main architecture with output $24\times 32\times 3$. Since we aim for a balance between resolution and achieving small errors on the test dataset, we therefore selected the model with the output resolution of $24\times 32\times 3$.
We also compare against training three separate U-net models, where each model is designed to solely predict a single force component, i.e., each individual model outputs a force distribution of shape $24\times 32\times 1$. Although these models collectively possess three times the number of parameters, their predictive accuracy is inferior to that of the main U-net w.r.t. the shear force estimates and only minorly improved w.r.t. normal force. This suggests that predicting all three force directions simultaneously can better capture correlations, especially between the shear forces, yielding more accurate predictions. Moreover, using one joint model also drastically improves the computational efficiency.
Lastly, when further increasing the output resolution to $48\times 64\times 3$, we still observe good performance; however, a slight decline compared to our main U-net.
Generalization to Novel Test Indenters
Table III: U-net Mean Absolute Error (MAE) on the Test Set containing new, previously unseen Indenters.
$\text{MAE}_{\text{GUF}}$ [N] $\text{MAE}_{\text{TF}}$ [N]
$f_{x}$ $0.0014\pm 0.0014$ $0.3010\pm 0.3777$
$f_{y}$ $0.0013\pm 0.0014$ $0.2193\pm 0.2837$
$f_{z}$ $0.0042\pm 0.0032$ $1.1862\pm 1.2998$
While the previous experiment evaluated our trained representation on the same indenters that were also used for training, this section now evaluates FEATS on $5$ new, previously unseen test indenters which were not included in the training data (cf. Fig. 2). We collected $2789$ samples with these new indenters, covering the same force range as for the other testing dataset (cf. Fig. 6). The results in Tab. III reveal that our trained network is indeed capable of generalizing to the previously unseen indenters. Considering the per-Grid Unit Forces (GUF), the errors only slightly increase by $0.0004\,\mathrm{N}$, $0.0003\,\mathrm{N}$, for the x- and y-direction, while the total force errors remain comparable to the previous experiment. In line with the higher range of the normal forces in this dataset, we also observe a slightly increased error of $0.0012\,\mathrm{N}$ compared to the evaluation on the known indenters. Since we only observe a small increase in the errors compared to the known indenters, we conclude that the representation generalizes well to these new contact configurations.
Generalization to Different Sensors
To explore the generalization capability of our FEATS method, we finally test it on images captured with a different GelSight Mini sensor of the same type. We again consider the previously unseen test indenters. As shown in Fig. 7, qualitatively, the predictions remain largely accurate. Quantitatively, we obtain mean total force errors of $0.24\,\mathrm{N}$, $0.27\,\mathrm{N}$, & $2.99\,\mathrm{N}$, for the $x$-, $y$-, and & $z$-direction, respectively. Similar to the results of the previous sections, we therefore again observe the largest increase in error for the normal force ($z$-direction), where also the data range is substantially larger. We hypothesize that the additional increase in error might be related with several factors, including differences in the alignment of the camera relative to the gel, and sensor-related factors such as image brightness, contrast, saturation and hue.
@@
Figure 7: Generalization to a different sensor. We evaluate our FEATS U-net on the previously unseen test indenters using a different GelSight Mini sensor to validate the model generalization capability. The first row displays the raw images captured by the sensor, and the subsequent rows show the components of the force distribution, in Newtons. We observe that the general outline of the force distribution is captured correctly despite the new sensor and the previously unseen indenters.